AIPMT

For an adiabatic expansion of a mono atomic perfect gas, the volume increases by 24% what is the percentage decrease in pressure

dasari sreecharan
8 Years agoGrade

1 Answer

Shubham Rai
8 Years ago
SINCE IN ADIABATIC EXPANSION
PVY= CONSTANT -(1)
LET INITIAL VOLUME BE V1 PRESSURE P1
AND FINAL VOLUME BE V2 AND PRESSURE P2
SINCE V2= V1+0.24V1= 1.24V1
AND Y= 5/3 FOR MONOATOMIC GAS
PUTTING VALUE OF V2 IN TERMS OF V1 IN (1) WE GET
P2= (1.24)5/3P1
PERCENTAGE IECREASE =  P2-P1/P1 × 100
=( 1.245/3-1) ×100 = 42.9 PERCENT
HENCE PRESSURE IS INCREASED NOT DECREASED